1. Suppose today is Dec 31st, 2019 and the U.S. Treasury notes have been issued with...

1.

Suppose today is Dec 31st, 2019 and the U.S. Treasury notes have been issued with a December 2024 maturity, $1000 face value, and a 2.2% coupon rate with semiannual coupons.

The first coupon payment will be paid on Jun. 30th 2020. If the yield to maturity is 3% today, then what is price of the U.S. Treasury notes today?

A. $1037.68

B. $970.06

C. $963.11

D. $837.93

2. Suppose a five-year, $1000 bond with semiannual coupons has a price of $936 and a yield to maturity of 6%. What is the bond's coupon rate as in APR?

______%

3.Here is the information of a Treasury bond (semi-annual coupon bond)

that matures in December 2020:

Maturity date    Coupon Rate     Bid                  Ask                  Ask Yield

12/31/2020 2.375 100.236 100.242 1.758

Which of the following statements is false?

Group of answer choices

A. The bond will pay a coupon of about $11.88 every six months.

B. If investors place a market buy order, they will pay $1002.42 for the bond. (For this question, don’t worry about the invoice price)

C. The bondholders will get an annual rate of return by 2.375% when holding the bond until maturity.

D. The bond’s yield is less than coupon rate, which suggests the bond is trading at a premium.

4.A semiannual coupon bond currently has an ask price of 115% of its $1,000 face value. The bond’s coupon rate is 6%. If the last coupon payment was made one month ago, what is the invoice price of the bond?
Answer #1

1.
=2.2%*1000/3%*(1-1/1.015^(5*2))+1000/1.015^(5*2)=963.1112618

2.
=(936-1000/1.03^10)*6%/1000*1/(1-1/1.03^10)=4.50%

3.
C. The bondholders will get an annual rate of return by 2.375% when holding the bond until maturity.

4.
=115%*1000+6%*1000/2*1/6=1155